OpenStudy (anonymous):

Find critical point(s) f(x)=(x+2)^2(x-1)

OpenStudy (anonymous):

I'm not sure what the first derivative will be...

OpenStudy (kamille):

\[f(x)=(x+2)^{2(x-1)}\] ?

OpenStudy (kamille):

or \[f(x)=(x+2)^{2}(x-1)\]

OpenStudy (anonymous):

The second way

OpenStudy (kamille):

do you know what critical point is?

OpenStudy (anonymous):

yes

OpenStudy (kamille):

so, firstly you need to find deretitive.do you know how?

OpenStudy (kamille):

\[f'(x)=2(x+2)(x-1)+(x+2)^{2}\] right or not?

OpenStudy (anonymous):

I think you are correct. I wasn't sure if I needed to simplify before deriving it.

OpenStudy (kamille):

Well, i think it doesnt matter when you simplify. NOW you need to simplify, can you do it for me?

OpenStudy (anonymous):

working on it...

OpenStudy (kamille):

\[f'(x)=2(x+2)(x-1)+(x+2)^{2}=2(x ^{2}-x+2x-2)+x ^{2}+4x+4=\]

OpenStudy (anonymous):

I got \[3x^2+6x+2\]

OpenStudy (kamille):

\[=2x ^{2}-2x+4x-4+x ^{2}+4x+4=3x ^{2}+6x\]

OpenStudy (kamille):

oh, we didnt get the same. maybe I have made a mistake!

OpenStudy (anonymous):

oops messed up on the 2

OpenStudy (anonymous):

that was my mistake!

OpenStudy (kamille):

oh,okay. so now do you know how to find critical point?

OpenStudy (anonymous):

So x=0 and x=-2 is what I got

OpenStudy (kamille):

well, \[3x ^{2}+6x=0\] \[x(3x+6)=0\] x=0 or x=-2  all points change a sign,so both of them are critical